OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (anonymous):

I just need an explanation on how factor them.

OpenStudy (anonymous):

I don't sure I think they are 10x2 − 7x − 3 8x2 − 11x + 3 because get ax^2+bx+c if a+b+c=0 we have ax^2+bx+c=a(x-x1)(x-x2) with x1=1 and x2=c/a